OpenStudy (anonymous):

Find the lim->2 of x^2-4x=4/x^2+3x-10

OpenStudy (anonymous):

what happens when you replace x by 2?

OpenStudy (anonymous):

i get 0 so need to use L'hopitals

OpenStudy (anonymous):

i assume this is \[\lim_{x\to 2}\frac{x^2-4x+4}{x^2+3x-10}\] and heck no, you do not need l'hopital, although it would work

OpenStudy (anonymous):

since 2 gives zero in both the numerator and denominator, you know you can factor each term s \[(x-2)(\text{something})\] so you can factor and cancel

OpenStudy (anonymous):

can u try it using l'hopitals i got 0/7

OpenStudy (anonymous):

\[\frac{(x-2)(x-2)}{(x-2)(x+5)}=\frac{x-2}{x+5}\] replace x by 2, get \[\frac{0}{7}=0\]

OpenStudy (anonymous):

or get \[\lim_{x\to 2}\frac{2x-4}{2x+3}\] gives the same thing

OpenStudy (anonymous):

second one was l'hopital, first one was factor and cancel your choice

OpenStudy (anonymous):

yea thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

wanna try another...just need confirmations